OpenStudy (anonymous):

(sqrt2-sqrt3)(sqrt2+sqrt3) is the answer 1?

OpenStudy (anonymous):

\[(\sqrt{2}-\sqrt{3})(\sqrt{2}+\sqrt{3})\]